OpenStudy (anonymous):

(a+2/a^2+a)/(a+2/a^3) simplify show work

OpenStudy (anonymous):

\[\frac{a+2 \over a^2 + a}{a+2 \over a^3}\] Simplify the complex fraction by multiplying the top and bottom of the outter most fraction by the least common-denominator of the inner fractions.

OpenStudy (anonymous):

So to be specific: What is the common denominator of \(a+2 \over a^2 + a\) and \(a+2 \over a^3\)

OpenStudy (anonymous):

Did I lose you jenni?

OpenStudy (anonymous):

polpak, wouldn't this approach simplify the process of finding a common denominator \[{{(a+2) \over (a^2+a)} \over {(a+2) \over a^3}} = {(a+2) \over (a^2+a)}* {a^3 \over (a+2)} \]

OpenStudy (anonymous):

You turned your fractions. I dislike turning fractions.

OpenStudy (anonymous):

I find it too easy to screw up.

OpenStudy (anonymous):

yea you lost me

OpenStudy (anonymous):

\[\frac{a+2 \over a^2 + a}{a+2 \over a^3} = \frac{a+2 \over a(a + 1)}{a+2 \over a^3}= \frac{a+2 \over a(a + 1)}{a+2 \over a^3} \cdot \frac{a^3(a+1)}{a^3(a+1)} = {a(a+2) \over (a+1)(a+2)} \]

OpenStudy (anonymous):

You have two inner fractions. One in the numerator, one in the denominator.

OpenStudy (anonymous):

You can find a least common denominator for those two fractions: lcd of \(a+2 \over a(a+1)\) and \(a+2\over a^3\) is \(a^3(a+1)\)

OpenStudy (anonymous):

If you multiply the numerator and denominator of your outer fraction by that lcd you will simplify the complex fraction.

OpenStudy (anonymous):

if i give a like can someone please show me how on here? http://www.dabbleboard.com/draw please

OpenStudy (anonymous):

Certainly.

OpenStudy (anonymous):

ok what is your email so i can ask you to join

OpenStudy (anonymous):

Lets use twiddla though. I find it easier to work with.

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

http://www.twiddla.com/576473

OpenStudy (anonymous):

i got the answer thanks Polpak :D